OpenStudy (lexluth0r):

On a coordinate grid, line AB has endpoint B at (24,16). The midpoint of line AB is P(4,-3). What is the coordinate of point A?

geerky42 (geerky42):

If you have points \((x_1,y_1)\) and \((x_2,y_2)\), then midpoint would be \(\left(\dfrac{x_1+x_2}{2},\dfrac{y_1+y_2}{2}\right)\) So here, we know that \(x_2 = 24,~y_2 = 16\), so from here, we can plug in \(\left(\dfrac{x_1+x_2}{2},\dfrac{y_1+y_2}{2}\right)\); \[4 = \dfrac{x_1+(24)}{2},-3 = \dfrac{y_1+(16)}{2}\] Can you solve \(x_1\) and \(y_1\) here?

OpenStudy (lexluth0r):

Hi, sorry Lol I was at school and the bell rang

OpenStudy (lexluth0r):

and lol, the bonnie gif

OpenStudy (lexluth0r):

Okay, so what you said, I tried the first time. It worked fine. The point should be (-16,-22), right? Then I wanted to check my work by finding the midpoint between that point and point B. That should match point P, right? Because it's the midpoint?

OpenStudy (lexluth0r):

@geerky42

geerky42 (geerky42):

Yeah. You got it right :)

OpenStudy (lexluth0r):

OKay, idk what I did wrong last time, but It did come out correct this time. Thanks for the help! C;

geerky42 (geerky42):

No problem :)